OpenStudy (anonymous):

MEDALS!!!! Jacob spends 60 minutes in the gym every day doing freehand exercises and running on the treadmill. He spends 30 minutes more running on the treadmill than doing freehand exercises. Part A: Write a pair of linear equations to show the relationship between the number of minutes Jacob does freehand exercises (y) and the number of minutes he runs on the treadmill(x). Part B: How much time does Jacob spend on doing freehand exercises? Part C: Is it possible for Jacob to have spent 40 minutes running on the treadmill? Explain your reasoning.

OpenStudy (anonymous):

I hate systems of equations, it's so annoying and I just need help working out these problems

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

here ill explain it

OpenStudy (anonymous):

Would it be… x = y + 30?

OpenStudy (anonymous):

what is your x and y?

OpenStudy (anonymous):

It says y is freehand exercises and x is running

OpenStudy (anonymous):

x being free hand and y being treadmill your system would be x+y=60 (because you spend 60 minutes doing both activities) y=x+30

OpenStudy (anonymous):

sorry according to your problem it would be x+y=60 x=y+30 this is part a

OpenStudy (anonymous):

Oh ok, that makes sense, I knew the second one just needed another one to make it systems of equations

OpenStudy (anonymous):

Part B: Plug in your value for x remember: x=y+30 so (y+30)+y=60 y+30+y=60 2y+30=60 2y=30 y=15 so 15 minutes doing freehand exercises

OpenStudy (anonymous):

Well then how would he fill the remaining 15 minutes?? Because if you plug that in then you only get 45

OpenStudy (anonymous):

I feel like that is wrong because 30 and 15 does not equal 60. let me check my math

OpenStudy (anonymous):

Yes I just saw that. Hang on

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

wait, no that is correct. Let me explain

OpenStudy (anonymous):

Jacob does not spend 30 minutes on the treadmill, he spends 30 minutes MORE than 15, which is 45. 15+45=60 y+x=60

OpenStudy (anonymous):

Oh ok that makes sense now

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

Do you need help with part c? You;re welcome

OpenStudy (anonymous):

I already have part C thank you